OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundreth. 5 = ln(x + 1)

OpenStudy (campbell_st):

raise each side as powers of e use the log/power law \[e^{\ln(a)} = a\] the equation is now \[e^5 = x + 1\] \[x = e^5 -1\] just evaluate

OpenStudy (anonymous):

Thanks [: